# Doomed Worlds I: No new evidence for orbital decay in a long-term survey of 43 ultra-hot Jupiters

### v1

@@ 6.3 CoRoT-2 b: no orbital decay
CoRoT-2 b, with d, was the second planet announced by the CoRoT space mission (Alonso et al., 2008), and it has one of the longest baselines in our survey, with data spanning 18 years (Table 3). Pre-discovery data using BEST show clear indications of transits two years before the CoRoT discovery, though only one of those light curves spanned both ingress and egress (Rauer et al., 2010). This planet has been frequently studied, since in addition to being one of the earliest known transiting planets, it orbits a highly active (7-20% of surface covered in spots, Guillot & Havel, 2011), rapidly rotating ( d), likely young (30-300 Myr) star (Gillon et al., 2010), and the planet itself has an inflated radius ( and ) that is a challenge to reproduce theoretically (Gillon et al., 2010). Previous work has found suggestions of non-linearity in the ephemeris: Öztürk & Erdem (2019) measured ms y, though with BIC values that favor a linear model, while Ivshina & Winn (2022), using additional data from TESS, found a much higher ms y. In both cases high scatter was seen in the timing residuals. Recently, Wang et al. (2023) found that the tiny error bars reported in Öztürk & Erdem (2019) strongly influenced the values for ; using more realistic errors they found a significant decreasing period with ms y, between the two previous estimates.
@@
We report seven new full or partial transits: three observed in 2008 (Adams, 2010), three observed in 2022-23 with the SAAO 40-inch telescope, and one in 2023 with Minerva-Australis. We have also fit 12 TESS transits from Sector 54, which was released since the publication of Ivshina & Winn (2022). Finally, we have re-fit the original CoRoT mission data, which observed 82 consecutive transits between 2007 May to October. We include in our timing analysis the reported transit midtimes for 36 light curves from the literature, as compiled by Ivshina & Winn (2022) after omitting duplicates and one composite time, plus 29 additional, non-duplicated light curves from the ETD, requiring midtime errors less than five minutes and DQ=1 (Poddaný et al., 2010). Such a large and diverse dataset is very useful for highlighting the challenges that arise with long-term timing analyses. We describe the case of CoRoT-2 b in detail to illustrate how more than one kind of error can coexist in the same dataset, and to illustrate the detective work that is required to both identify and fix the errors.
@@
Our initial examination of the system added the seven new ground-based transits plus TESS Sector 54 to the midtimes collected by Ivshina & Winn (2022), but did not re-analyze the CoRoT mission times. Using these data, CoRoT-2 b had the second strongest signature for a non-linear orbital period in our sample (after WASP-12 b), with a 3.1- detection of a decreasing period and a value of 383. The initial best fit rate of decay was calculated to be ms y. This value was consistent with what Öztürk & Erdem (2019) reported and would have implied a very low value for , and thus we subjected the data to additional scrutiny.
@@
Underestimated errors. The second issue involves the published times for the CoRoT mission data from Öztürk & Erdem (2019). The CoRoT mission discovery dataset contains 82 consecutive, high-quality transits at the very beginning of the time series, and are highly influential on the observed value. However, the errors on individual transit midtimes reported in Ivshina & Winn (2022) and originally published in Öztürk & Erdem (2019) were much smaller than the scatter between midtime residuals (O-C values), indicating that the errors were probably underestimated, as noted by Wang et al. (2023). Indeed, Figure 1 in Öztürk & Erdem (2019) shows adjusted transit midtime errors that are times higher, though their table (perhaps inadvertently) published the smaller errors. Low error bars on 79 precise points at an early epoch have a huge impact on the value, so correcting these errors was of critical importance.
@@
Strong unexplained trend. The most serious issue was a trend that was apparent by eye in the 150 days of CoRoT mission data from Öztürk & Erdem (2019). This trend would have implied a faster rate of decay than that of WASP-12 b and . In contrast to WASP-12 b, for which the star may be slightly evolved (see Section 8), if anything the evidence points toward CoRoT-2 being a young star, with corresponding high values expected for . Removing the CoRoT points entirely caused the value to drop near zero, so it was clear that any apparent detection of orbital period decrease rested on the timing of these transits.
To get to the bottom of both (2) and (3), we downloaded the CoRoT mission data and refit all 82 transits (see Section 5). Using our new times, which are shown in Table 5, we re-ran the timing analysis and found that , indicating no orbital decay. Our timing results are show in Figure 5. The differences between our values for the CoRoT mission times and the midtimes reported by Öztürk & Erdem (2019) range from 200-500 s and follow quite closely the expected barycentric correction at each midtime. Thus we hypothesize that Öztürk & Erdem (2019) used JD times in their light curves instead of either HJD/TT (in the older version of the CoRoT pipeline that they worked with) or BJD/TDB (in the Legacy version used in this work). Unfortunately, this error was then propagated to Ivshina & Winn (2022), which labeled the data as being in BJD/TDB.
@@ 6.4 WASP-121 b: possible period increase heavily dependent on single composite time at earliest epoch
WASP-121 b has d and was discovered by Delrez et al. (2016). We report three new light curves from 2022-2023 with the SAAO 40-inch (Table 1), and use an additional 59 times from the literature and 6 times from the ETD. The first point in the sequence is a composite of nine high-quality transits taken between 9 Dec 2013 and 29 Dec 2014 (Delrez et al., 2016). Note that all but five of the literature times are from two TESS sectors, making this a very sparsely observed target, with large gaps in timing coverage. For that reason, we did not remove the initial composite time, since it would have removed years of timing baseline. Fitting those times individually would provide an excellent constraint on the timing, but unfortunately these data have not been published outside of a figure in the discovery paper.
Our timing analysis is shown in Figure 6. Not surprisingly, the value for is highly dependent on the midtime from Delrez et al. (2016), as shown by the omit-one test; removing it causes the effect to entirely vanish as drops below 0. However, this system is also the only other system (besides WASP-12 b) that does not see its value evaporate when the error bars are rescaled upward (see Figure 2), retaining the second-highest rescaled (all data, rescaled errors). This indicates that the system has reasonable error bars on most data, but needs additional epochs of observations, both through additional new light curves and a reanalysis, if possible, of the oldest data as individual transit midtimes.
@@ 6.5 WASP-46 b: evidence for period increase is murky
WASP-46 b, with d, was discovered by Anderson et al. (2012) and is shown in Figure 7. We observed three new transits in 2023 with the SAAO 40-inch telescope (Table 1) and fit 20 transits from the latest TESS Sector (Sector 67). In the timing analysis, we used 86 transit midtimes from Ivshina & Winn (2022) (excluding one composite time), which included two sectors of TESS data. We also used 28 non-duplicated light curves from the ETD with DQ=1 or 2 and errors under 5 minutes.
Several previous timing analyses of WASP-46 b exist, including Ciceri et al. (2016); Moyano et al. (2017); Petrucci et al. (2018); Davoudi et al. (2021). Since each of these papers also included different subsets of ETD transits, care had to be taken to avoid duplication of data points. No evidence for timing change was seen in Petrucci et al. (2018), though an increase in period was noted by Davoudi et al. (2021). We see a preference for a quadratic model with positive curvature, with =109 and a 2.6 detection of ms yr. However, it is important to note that the data before about 2015 has very high scatter in the transit-timing residuals, and seven points have been flagged as outliers which, if omitted individually, cause the value for to range from 32 to 149 (see Figure 7). No obvious signs of timing conversion errors were found on investigating the source of those early midtimes. We also note that when the errors are rescaled, the value of plummets to 2 – still a slight preference for a non-linear model and among the highest values reported for rescaled values (see Figure 2), but by no means a smoking gun. To resolve this murkiness, we recommend WASP-46 b for further analysis and regular transit observations.
@@ 6.6 TrES-1 b: period potentially decreasing, but orbital decay unlikely
TrES-1 b, with a period of d, was one of the first transiting planets discovered (Alonso et al., 2004) and has a very long baseline of observations, as well as prior tentative claims of period decrease (Ivshina & Winn, 2022; Hagey et al., 2022). Two transits of TrES-1b were taken with MORIS on the IRTF in 2009, and we have also fit the two most recent sectors of TESS data (Sectors 53 and 54). We used 61 previously published midtimes from Ivshina & Winn (2022) and another 49 midtimes from the ETD.
@@
In contrast with other systems, no obvious errors were identified in the literature times for TrES-1 b, and none of the midtimes were significant in our omit-one test. We found and a 3.4 detection of ms yr (see Figure 8). This is intermediate to, and consistent with, the value of ms yr that Ivshina & Winn (2022) found, and the value of ms yr from Hagey et al. (2022), and less significant than either. However, assuming the apparent period decrease holds up to further observations, the explanation will likely be something other than orbital decay, since the inferred is five orders of magnitude lower than the star’s predicted (Weinberg et al., 2024). Even the rapidly decaying Kepler-1658 b, which orbits a subgiant star, has an estimated value of , and there are no indications that the star TrES-1 has left the main sequence.
@@
A rapid ephemeris shift could result from perturbations from a massive companion. A stellar companion, TrES-1 B, is located 13.2 arcsec away, and is 8.5 magnitudes fainter, though it is not clear if a star that distant could produce a perturbation with the magnitude and minimum oscillation period implied by this detection, or if some other, as yet undetected body would need to be responsible. Exploring the full range of dynamical scenarios that could lead to such a rapid are beyond the scope of this work, but based on our analysis the signature for a non-linear period is promising, and we recommend TrES-1 b for continued monitoring.
6.7 Kepler-1658 b: existing data has promising
@@
We applied our analysis to another promising decay candidate, Kepler-1658b, a d planet that orbits an evolved host star (, ). The orbital period decrease for this planet was measured as ms yr by Vissapragada et al. (2022). The evolved host star is probably key to explaining the rapid rate of decay: although Weinberg et al. (2024) found that g-mode dissipation is too weak to explain it, Vissapragada et al. (2022) argued that inertial wave dissipation, combined with rapid stellar rotation, could.On the other hand, recent results from Barker et al. (2024) finds that the time the star spends with may be too short, potentially requiring non-synchronous rotation if tidal dissipation is to explain the rate of orbital decrease. An alternative explanation could be tidally-induced apsidal precession (Barker et al., 2024).
@@
Although we have no new transit observations for this planet due to its very shallow transit depth (0.1%), we applied our analysis pipeline to the mid-transit times reported in Vissapragada et al. (2022) as described above. We found ms yr, with , in agreement with Vissapragada et al. (2022). The high scatter in the photometry compared to the depth of individual transits meant that it was necessary to stack all transits in each Kepler quarter or TESS sector to produce a single time per quarter/sector. In addition, three ground-based observations were made using the Palomar 5.1-m (Vissapragada et al., 2022); one of these points triggers our omit-one test, and without it the value drops in half. Since this point has low errors and occurs near the end of the sequence, it is not surprising that it holds so much weight; this does not necessarily indicate a problem with the data but does point to the need for additional observations. We note that the nominal and rescaled values for are nearly identical (49 and 50, respectively), indicating that the reported errors are likely reliably measured. The extremely rapid rate of decay for this system means that confirmation could be possible with another year or two of observations.
@@
| b. HAT-P-23 theoretical calculations: if use , and . | ||||||||||
| c. HAT-P-23 theoretical calculations: if use , and . | ||||||||||
@@ 6.8 Data for marginal systems ()
We chose somewhat arbitrarily to select a handful of the most promising candidates to focus on in this paper. It is also a useful discriminant of where values as calculated with real data and error bars start to become robust enough to have more confidence that they are measuring something other than noise. Many of the systems with possible claimed detections listed in the introduction were presented at much smaller values. We will now briefly discuss our analysis of the 37 planets that fall below which have not been described above, as shown in Figure 2. Our preliminary values for and the best-fit quadratic model parameters are listed in Table 7. We add an important caveat that unlike with the 6 main targets of this paper, for these 37 planets we have not done a deep dive into the sources of the literature times used. As we have demonstrated for other systems above, it is quite possible that there may be additional problems lurking in the data for one or more of these systems.
@@
For each of these 37 planets, we used literature times from Ivshina & Winn (2022) and from the ETD. We removed duplicate entries, and have limited the ETD data to DQ=1 or 2 for most targets. (For WASP-163b, which has very few ETD points, we allowed all DQ=1-5, while for Qatar-1b, TrES-3b, and WASP-43b, which have many transits, we used only DQ=1.) We have not searched for additional transit midtimes that were not in Ivshina & Winn (2022), including systems for which there is known Kepler or K2 data. We have, however, analyzed additional sectors of TESS data where available. For systems that are recently discovered or have very few published literature times, we refit some literature light curves that had not previously had individual fits, using data provided in private correspondence with the authors of those works. These systems are: HATS-70 b (6 light curves, Zhou et al., 2019); NGTS-1 b (1 light curve and 1 NGTS composite light curve, Bayliss et al., 2018); Qatar-10 b Alsubai et al. (7 light curves, 2019); WASP-43 b Fox & Wiegert (3 light curves, 2022). These analyses will be part of forthcoming publications similar to this one, and the values presented in Table 7 should be considered preliminary results.
@@
For all systems with nominal , few retained a positive value for when we rescaled the error bars to determine the impact of underestimated errors on (see discussion of test in Section 5.4). Of the four high () planets, all still have positive rescaled values, as noted above. But among the four intermediate planets (), only two, WASP-164 b and TrES-3 b, still have positive rescaled values; and among the 8 planets with slightly positive nominal values (), just one, TOI-2046 b, is still positive in the rescaled test. Two other planets, TOI-564 b and HIP 65A b, had small negative nominal values that switched to small positive rescaled values. The last three planets are all sparsely sampled with large errors on individual transits and only 4-5 years of baseline, precisely the kinds of systems where swings in are common (Jackson et al., 2023), and more data are needed before the true patterns of these systems will become clear. We interpret the shifts in to be suggestive that underestimated errors alone may explain most of the marginal detections of decreasing orbital period, both here and in the literature.
@@
Of the four planets in Figure 2 with intermediate values of , TrES-3 b, WASP-52 b, WASP-164 b, and WASP-32 b, all have decreasing periods, though none have significant values of (). Of these, two have previous claimed detections: (1) A possible decreasing period for TrES-3 b has been noted at similar marginal levels by other works (Mannaday et al., 2022; Hagey et al., 2022). Given that TrES-3 b is a long-studied planet around a bright star with particularly deep transits, hundreds of diverse light curves have been recorded by amateurs and professionals alike (with over 600 light curves listed in the ETD alone), making the chance of there being undetected mistakes in the literature high. A thorough reanalysis of the available record needs to be undertaken before any claims for or against period change could be credibly believed for TrES-3 b. (2) Sun et al. (2023) noted possible non-linearity for WASP-32 b, though that work preferred an apsidal precession model to the quadratic orbital decay model. Neither their value for nor ours is significant (both around 2 ), and the system is relatively sparsely observed. Additional observations are required to assess this claim.
7 Preliminary population analysis: where are the decaying Ultra-hot Jupiters?

### v2

@@ 6.3 CoRoT-2 b: no orbital decay
CoRoT-2 b, with d, was the second planet announced by the CoRoT space mission (Alonso et al., 2008), and it has one of the longest baselines in our survey, with data spanning 18 years (Table 3). Prediscovery data using BEST show clear indications of transits two years before the CoRoT discovery, though only one of those light curves spanned both ingress and egress (Rauer et al., 2010). This planet has been frequently studied, since in addition to being one of the earliest known transiting planets, it orbits a highly active (7-20% of the surface covered in spots, Guillot & Havel, 2011), rapidly rotating ( d), likely young (30-300 Myr) star (Gillon et al., 2010), and the planet itself has an inflated radius ( and ) that is a challenge to reproduce theoretically (Gillon et al., 2010). Previous work has found suggestions of nonlinearity in the ephemeris: Öztürk & Erdem (2019) measured ms y-1, though with BIC values that favor a linear model, while Ivshina & Winn (2022), using additional data from TESS, found a much higher ms y-1. In both cases high scatter was seen in the timing residuals. Recently, Wang et al. (2024) found that the tiny error bars reported in Öztürk & Erdem (2019) strongly influenced the values for ; using more realistic errors they found a significantly decreasing period with ms y-1, between the two previous estimates.
@@
We report seven new full or partial transits: three observed in 2008 (Adams, 2010), three observed in 2022-2023 with the SAAO 40 inch telescope, and one observed in 2023 with Minerva-Australis. We have also fit 12 TESS transits from Sector 54, which was released after the publication of Ivshina & Winn (2022). Finally, we have refit the original CoRoT mission data, which observed 82 consecutive transits between 2007 May and October. We include in our timing analysis the reported transit midtimes for 36 light curves from the literature, as compiled by Ivshina & Winn (2022) after omitting duplicates and one composite time, plus 29 additional, nonduplicated light curves from the ETD, requiring midtime errors less than five minutes and DQ=1 (Poddaný et al., 2010). Such a large and diverse dataset is very useful for highlighting the challenges that arise with long-term timing analyses. We describe the case of CoRoT-2 b in detail to illustrate how more than one kind of error can coexist in the same data set and to illustrate the detective work that is required to both identify and fix the errors.
@@
Our initial examination of the system added the seven new ground-based transits plus TESS Sector 54 to the midtimes collected by Ivshina & Winn (2022), but did not reanalyze the CoRoT mission times. Using these data, CoRoT-2 b had the second strongest signature for a non-linear orbital period in our sample (after WASP-12 b), with a 3.1- detection of a decreasing period and a value of 383. The initial best-fit rate of decay was calculated to be ms y-1. This value was consistent with what Öztürk & Erdem (2019) reported and would have implied a very low value for , and thus we subjected the data to additional scrutiny.
@@
Underestimated errors. The second issue involves the published times for the CoRoT mission data from Öztürk & Erdem (2019). The CoRoT mission discovery dataset contains 82 consecutive, high-quality transits at the very beginning of the time series, and are highly influential on the observed value. However, the errors on individual transit midtimes reported in Ivshina & Winn (2022) and originally published in Öztürk & Erdem (2019) were much smaller than the scatter between midtime residuals (O-C values), indicating that the errors were probably underestimated, as noted by 2024ApJS..270...14. Indeed, Figure 1 in Öztürk & Erdem (2019) shows adjusted transit midtime errors that are times higher, though their table (perhaps inadvertently) published the smaller errors. Low error bars on 79 precise points at an early epoch have a huge impact on the value, so correcting these errors was of critical importance.
@@
Strong unexplained trend. The most serious issue was a trend that was apparent by eye in the 150 days of CoRoT mission data from Öztürk & Erdem (2019). This trend would have implied a faster rate of decay than that of WASP-12 b and . In contrast to WASP-12 b, for which the star may be slightly evolved (see Section 8), if anything the evidence points toward CoRoT-2 being a young star, with corresponding high values expected for . Removing the CoRoT points entirely caused the value to drop near 0, so it was clear that any apparent detection of orbital period decrease rested on the timing of these transits.
To get to the bottom of both (2) and (3), we downloaded the CoRoT mission data and refit all 82 transits (see Section 5). Using our new times, which are shown in Table 5, we re-ran the timing analysis and found that , indicating no orbital decay. Our timing results are show in Figure 6. The differences between our values for the CoRoT mission times and the midtimes reported by Öztürk & Erdem (2019) range from 200 to 500 s and follow quite closely the expected barycentric correction at each midtime. Thus we hypothesize that Öztürk & Erdem (2019) used JD times in their light curves instead of either HJD/TT (in the older version of the CoRoT pipeline that they worked with) or BJD/TDB (in the Legacy version used in this work). Unfortunately, this error was then propagated to Ivshina & Winn (2022), who labeled the data as being in BJD/TDB.
@@ 6.4 WASP-121 b: Possible Period Increase Heavily Dependent on Single Composite Time at Earliest Epoch
WASP-121 b has d and was discovered by Delrez et al. (2016). We report three new light curves from 2022 to 2023 with the SAAO 40-inch (Table 1) and use an additional 59 times from the literature and 6 times from the ETD. The first point in the sequence is a composite of nine high-quality transits taken between 9 December 2013 and 29 December 2014 (Delrez et al., 2016). Note that all but five of the literature times are from two TESS sectors, making this a very sparsely observed target, with large gaps in timing coverage. For that reason, we did not remove the initial composite time, since it would have removed years of timing baseline. Fitting those times individually would provide an excellent constraint on the timing, but unfortunately these data have not been published outside of a figure in the discovery paper.
Our timing analysis is shown in Figure 7. Not surprisingly, the value for is highly dependent on the midtime from Delrez et al. (2016), as shown by the omit-one test; removing it causes the effect to entirely vanish as drops below 0. However, this system is also the only other system (besides WASP-12 b) that does not see its value evaporate when the error bars are rescaled upward (see Figure 2), retaining the second-highest rescaled (all data, rescaled errors). This indicates that the system has reasonable error bars on most data but needs additional epochs of observations, through both additional new light curves and a reanalysis, if possible, of the oldest data as individual transit midtimes.
@@ 6.5 WASP-46 b: Evidence for Period Increase Is Murky
WASP-46 b, with d, was discovered by Anderson et al. (2012) and is shown in Figure 8. We observed three new transits in 2023 with the SAAO 40 inch telescope (Table 1) and fit 20 transits from the latest TESS Sector (Sector 67). In the timing analysis, we used 86 transit midtimes from Ivshina & Winn (2022) excluding one composite time, which included two sectors of TESS data. We also used 28 nonduplicated light curves from the ETD with DQ=1 or 2 and errors under 5 minutes.
Several previous timing analyses of WASP-46 b exist, including Ciceri et al. (2016); Moyano et al. (2017); Petrucci et al. (2018); Davoudi et al. (2021). Since each of these papers also included different subsets of ETD transits, care had to be taken to avoid duplication of data points. No evidence for timing change was seen in Petrucci et al. (2018), though an increase in period was noted by Davoudi et al. (2021). We see a preference for a quadratic model with positive curvature, with =109 and a 2.6 detection of ms yr-1. However, it is important to note that the data before about 2015 have very high scatter in the transit timing residuals, and seven points have been flagged as outliers that, if omitted individually, cause the value for to range from 32 to 149 (see Figure 8). No obvious signs of timing conversion errors were found on investigating the source of those early midtimes. We also note that when the errors are rescaled, the value of plummets to 2 – still a slight preference for a nonlinear model and among the highest values reported for rescaled values (see Figure 2), but by no means a smoking gun. To resolve this murkiness, we recommend WASP-46 b for further analysis and regular transit observations.
@@ 6.6 TrES-1 b: Period Potentially Decreasing, But Orbital Decay Unlikely
TrES-1 b, with a period of d, was one of the first transiting planets discovered (Alonso et al., 2004) and has a very long baseline of observations, as well as prior tentative claims of period decrease (Hagey et al., 2022; Ivshina & Winn, 2022). Two transits of TrES-1 b were taken with MORIS on the IRTF in 2009, and we have also fit the two most recent sectors of TESS data (Sectors 53 and 54). We used 61 previously published midtimes from Ivshina & Winn (2022) and another 49 midtimes from the ETD.
@@
In contrast with other systems, no obvious errors were identified in the literature times for TrES-1 b, and none of the midtimes were significant in our omit-one test. We found and a 3.4 detection of ms yr-1 (see Figure 9). This is intermediate to, and consistent with, the value of ms yr-1 that Ivshina & Winn (2022) found, and the value of ms yr-1 from Hagey et al. (2022), and less significant than either. However, assuming the apparent period decrease holds up to further observations, the explanation will likely be something other than orbital decay, since the inferred is five orders of magnitude lower than the star’s predicted (Weinberg et al., 2024). Even the rapidly decaying Kepler-1658 b, which orbits a subgiant star, has an estimated value of , and there are no indications that the star TrES-1 has left the main sequence.
@@
A rapid ephemeris shift could result from perturbations from a massive companion. A stellar companion is located 13.2 arcsec away and is 8.5 magnitudes fainter, though it is not clear if a star that distant could produce a perturbation with the magnitude and minimum oscillation period implied by this detection or if some other as-yet-undetected body would need to be responsible. Exploring the full range of dynamical scenarios that could lead to such a rapid is beyond the scope of this work, but based on our analysis the signature for a nonlinear period is promising, and we recommend TrES-1 b for continued monitoring.
6.7 Kepler-1658 b: Existing Data Have Promising
@@
We applied our analysis to another promising decay candidate, Kepler-1658b, a d planet that orbits an evolved host star (, ). The orbital period decrease for this planet was measured as ms yr-1 by Vissapragada et al. (2022). The evolved host star is probably key to explaining the rapid rate of decay: although Weinberg et al. (2024) found that g-mode dissipation is too weak to explain it, Vissapragada et al. (2022) argued that inertial wave dissipation, combined with rapid stellar rotation, could.On the other hand, recent results from Barker et al. (2024) find that the time the star spends with may be too short, potentially requiring nonsynchronous rotation if tidal dissipation is to explain the rate of orbital decrease. An alternative explanation could be tidally induced apsidal precession (Barker et al., 2024).
@@
Although we have no new transit observations for this planet due to its very shallow transit depth (0.1%), we applied our analysis pipeline to the midtransit times reported in Vissapragada et al. (2022) as described above. We found ms yr-1, with , in agreement with Vissapragada et al. (2022); see Figure 10. The high scatter in the photometry compared to the depth of individual transits meant that it was necessary to stack all transits in each Kepler quarter or TESS sector to produce a single time per quarter/sector. In addition, three ground-based observations were made using the Palomar 5.1-m (Vissapragada et al., 2022); one of these points triggers our omit-one test, and without it the value drops by half. Since this point has low errors and occurs near the end of the sequence, it is not surprising that it holds so much weight; this does not necessarily indicate a problem with the data but does point to the need for additional observations. We note that the nominal and rescaled values for are nearly identical (49 and 50, respectively), indicating that the reported errors are likely reliably measured. The extremely rapid rate of decay for this system means that confirmation could be possible with another year or two of observations.
@@
| b. HAT-P-23 theoretical calculations if using , and . | ||||||||||
| c. HAT-P-23 theoretical calculations if using , and . | ||||||||||
@@ 6.8 Data for Marginal Systems ()
We chose somewhat arbitrarily to select a handful of the most promising candidates to focus on in this paper. It is also a useful discriminant of where values as calculated with real data and error bars start to become robust enough to have more confidence that they are measuring something other than noise. Many of the systems with possible claimed detections listed in the Introduction were presented at much smaller values. We will now briefly discuss our analysis of the 37 planets that fall below that have not been described above, as shown in Figure 2. Our preliminary values for and the best-fit quadratic model parameters are listed in Table 7. We add an important caveat that unlike with the 6 main targets of this paper, for these 37 planets we have not done a deep dive into the sources of the literature times used. As we have demonstrated for other systems above, it is quite possible that there may be additional problems lurking in the data for one or more of these systems.
@@
For each of these 37 planets, we used literature times from Ivshina & Winn (2022) and the ETD. We removed duplicate entries and have limited the ETD data to DQ=1 or 2 for most targets. (For WASP-163 b, which has very few ETD points, we allowed all DQ=1-5, while for Qatar-1 b, TrES-3 b, and WASP-43 b, which have many transits, we used only DQ=1.) We have not searched for additional transit midtimes that were not in Ivshina & Winn (2022), including systems for which there is known Kepler or K2 data. We have, however, analyzed additional sectors of TESS data where available. For systems that are recently discovered or have very few published literature times, we refit some literature light curves that had not previously had individual fits, using data provided in private correspondence with the authors of those works. These systems are: HATS-70 b (6 light curves, Zhou et al., 2019); NGTS-1 b (1 light curve and 1 NGTS composite light curve, Bayliss et al., 2018); Qatar-10 b Alsubai et al. (7 light curves, 2019); WASP-43 b Fox & Wiegert (3 light curves, 2022). These analyses will be part of forthcoming publications similar to this one, and the values presented in Table 7 should be considered preliminary results.
@@
For all systems with nominal , few retained a positive value for when we rescaled the error bars to determine the impact of underestimated errors on (see discussion of test in Section 5.4). Of the four high- () planets, all still have positive rescaled values, as noted above. But among the four intermediate planets (), only two, WASP-164 b and TrES-3 b, still have positive rescaled values, and among the 8 planets with slightly positive nominal values (), just one, TOI-2046 b, is still positive in the rescaled test. Two other planets, TOI-564 b and HIP 65A b, had small negative nominal values that switched to small positive rescaled values. The last three planets are all sparsely sampled with large errors on individual transits and only 4-5 years of baseline, precisely the kinds of systems where swings in are common (Jackson et al., 2023), and more data are needed before the true patterns of these systems will become clear. We interpret the shifts in to be suggestive that underestimated errors alone may explain most of the marginal detections of decreasing orbital period, both here and in the literature.
@@
Of the four planets in Figure 2 with intermediate values of , TrES-3 b, WASP-52 b, WASP-164 b, and WASP-32 b, all have decreasing periods, though none have significant values of (). Of these, two have previous claimed detections: (1) A possible decreasing period for TrES-3 b has been noted at similar marginal levels by other works (Hagey et al., 2022; Mannaday et al., 2022). Given that TrES-3 b is a long-studied planet around a bright star with particularly deep transits, hundreds of diverse light curves have been recorded by amateurs and professionals alike (with over 600 light curves listed in the ETD alone), making the chance of there being undetected mistakes in the literature high. A thorough reanalysis of the available record needs to be undertaken before any claims for or against period change could be credibly believed for TrES-3 b. (2) Sun et al. (2023) noted possible nonlinearity for WASP-32 b, though that work preferred an apsidal precession model to the quadratic orbital decay model. Neither their value for nor ours is significant (both around 2 ), and the system is relatively sparsely observed. Additional observations are required to assess this claim.
7 Preliminary Population Analysis: Where Are the Decaying UHJs?